OpenStudy (anonymous):

minimize Q=x^2+2y^2, where x+y=3

OpenStudy (anonymous):

The first step is to rewrite x+y=3, as y=3-x

OpenStudy (anonymous):

kk i got that part

OpenStudy (anonymous):

Now we can replace the y in the equation with our newfound value for y, namely y=3-x. : Q=x^2+2(3-y)^2 =x^2+18-12x+2x^2 =3x^2+18-12x

OpenStudy (anonymous):

Then we have to determine where the minimum of this parabola is

OpenStudy (anonymous):

We can do this simply by finding the vertex

OpenStudy (anonymous):

do you know how to find the vertex?

OpenStudy (anonymous):

yes get the derivative and set it = to 0

OpenStudy (anonymous):

x=2 y=1 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you took a misstep somewhere

OpenStudy (anonymous):

Q'(x)= 6x-12

OpenStudy (anonymous):

the y value should be 6

OpenStudy (anonymous):

no x+y=3 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6x-12=0 6x=12 x=2

OpenStudy (anonymous):

plug it into the original equation

OpenStudy (anonymous):

I did and q=6

OpenStudy (anonymous):

not y :)

OpenStudy (anonymous):

if x=2 then 2+1=3 y=1 x+y=3

OpenStudy (amistre64):

another method is a lagrange multiplier :)

OpenStudy (anonymous):

we dont even need anything else, all we need to do is find the vertex

OpenStudy (anonymous):

no method was specified

OpenStudy (amistre64):

Q=x^2+2y^2; constrained by: x+y=3 F(x,y,L) = x^2+2y^2 -L(x+y-3) Fx = 2x - L 2x - L = 0 x = L/2 Fy = 4y - L 4y - L = 0 y = L/4 FL = x+y-3 x+y-3 = 0 L/2+L/3-3 = 0 L(1/2+1/3) = 3 L(5/6) = 3 L = 18/5 x = L/2 = 9/5 y = L/4 =9/20 if i did it right

OpenStudy (anonymous):

the correct answer was Q=6 x=2 y=1

OpenStudy (anonymous):

from the back of the book

OpenStudy (amistre64):

hmmm .... guess my lagranging needs more work; either that or i found the maximum :)

OpenStudy (anonymous):

lol it was a u shaped parabola

OpenStudy (anonymous):

it is: Q=x^2+2(3-x)^2 q=3x^2-12x+18 q'(x)=6x-12 x=2 2+y=3 y=1

OpenStudy (anonymous):

(2)^2+ 2(1)^2=6

OpenStudy (anonymous):

q=6